Hi I need some help with this problem: Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution 9, only solution.
thanks
Charmaine

Found 2 solutions by stanbon, rwm:
Answer by stanbon(75887)

Hi I need some help with this problem: Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution 9, only solution.
thanks
----
Since 9 is the only solution and the equation
is quadratic, 9 must have multiplicity 2.
---
Equation:
y = (x-9)^2
y = x^2-18x+81
===================
cheers,
Stan H.

Answer by rwm(914)

Under what conditions does a quadratic equation have only one solution?
see http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.134718.html
for an explanation
(x-9)^2=0
x^2-18x+81=0
there will be two solutions that are equal or only one unique solution.